[ A motorcyclist makes a trip of 500 km. If he had increased his speed 10 kph, he would have covered 600 km in the same time. What was his original speed? ]
Answer by TimothyLamb(4379)

---
s = original speed
---
s = d / t
t = d / s
---
t = 500/s
t = 600/(s + 10)
---
500/s = 600/(s + 10)
500/s - 600/(s + 10) = 0
500(s + 10)/s(s + 10) - 600s/s(s + 10) = 0
(500s + 5000)/s(s + 10) - 600s/s(s + 10) = 0
500s + 5000 - 600s = 0
5000 - 100s = 0
100s = 5000
s = 5000/100
---
solution:
s = 50 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php